An antifreeze solution is made by mixing ethylene glycol (ρ = 1116 kg/m3) with water. Suppose the specific gravity of such a solution is 1.05. Assuming that the total volume of the solution is the sum of its parts, determine the volume percentage of ethylene glycol in the solution.
Answer #1

here,

density of the ethylene glycol , pe = 1116 kg/m^3

specific gravity of the solution is 1.05

density of solution , p = 1050 kg/m^3

let the volume of the solution be 1 m^3

let the volume of ethyle alchoal be x

mw + me = ms

Vw * pw + Ve *pe = Vs*ps

(1 - x) *1000 + x*1116 = 1 *1050

x = 0.43103 m^3

the volume percentage of ethylene glycol in the solution = 0.43103/1 * 100

= 43.1 %

the volume percentage of ethylene glycol in the solution is 43.1 %